OpenStudy (anonymous):

A bag contains 8 red and 4 white balls. If A, B, C draw them in succession, wat is the probability that C will get a white ball?

OpenStudy (anonymous):

Can anyone also suggest a good book for probability?

OpenStudy (asnaseer):

first - can you write down the different ways in which C can end up drawing a white ball? e.g. A could draw a red ball, then B could draw a red ball, then C could draw a white ball

OpenStudy (anonymous):

yup.... then?

OpenStudy (asnaseer):

if you write those down first then, for each row, you work out the probability of that scenario happing. then you add up the probabilities across all rows.

OpenStudy (asnaseer):

*happening

OpenStudy (anonymous):

Is it like this\[[(8*7*4)+(4*8*3)+(8*4*3)+(4*3*2)]/(12*11*10)\]

OpenStudy (asnaseer):

yes - that looks correct

OpenStudy (anonymous):

but its not giving me the correct answer...answer should be 7/33. is there any othere concept?

OpenStudy (asnaseer):

not that I can think of.

OpenStudy (asnaseer):

unless the question really means: A, B then C will draw a ball each. and this will continue until no balls are left. what is the probability that C has drawn a white ball?

OpenStudy (anonymous):

yup...this is it...but then it gets even more complex....how do i get it then

OpenStudy (asnaseer):

you can then use the same method for each round of drawing

OpenStudy (asnaseer):

there must be a way of expressing this in a concise manner

OpenStudy (asnaseer):

maybe if you consider the case where C does not draw any white ball at all, and then subtract that from 1 in the end?

OpenStudy (anonymous):

that will complicate the sum even further....pls elaborate..

OpenStudy (asnaseer):

thinking...

OpenStudy (anonymous):

take your time....i want to learn..

OpenStudy (asnaseer):

ok - first thought is as follows: probability that C does NOT draw a white in first round is 8/12 now we need to consider what combination of balls could be left after the first draw.

OpenStudy (anonymous):

means in the second trial there only two white balls left...

OpenStudy (asnaseer):

so I get after first round there are either: a) 7 Red, 2 White b) 6 Red, 3 White c) 5 Red, 4 White

OpenStudy (asnaseer):

now we need to consider case (a) to (c) for the second round and work out the probability of C NOT drawing a white ball

OpenStudy (anonymous):

getting really jammed...

OpenStudy (asnaseer):

which I /think/ is as follows: a) 7 Red, 2 White: probability of C NOT drawing a white = 7/9 b) 6 Red, 3 White: probability of C NOT drawing a white = 6/9 c) 5 Red, 4 White: probability of C NOT drawing a white = 5/9

OpenStudy (anonymous):

do we add all this.....or multiply all of them...

OpenStudy (asnaseer):

I /think/ we will need to do this: (8/12)( (7/9) + (6/9) + (5/9) )

OpenStudy (asnaseer):

no - that cannot be right

OpenStudy (anonymous):

the answer is greater than one...not possible...

OpenStudy (asnaseer):

hmmm - this case is trickier than I thought - let me think about it and I'll get back to you (unless someone else solves this before I get back)

OpenStudy (asnaseer):

?

OpenStudy (anonymous):

okkk....@asnaseer

OpenStudy (anonymous):

@waterineyes ..can u help

OpenStudy (anonymous):

Is the above that you have posted question??

OpenStudy (anonymous):

yup...it will kind of u if u try with the problem..

OpenStudy (anonymous):

Let me think..

OpenStudy (anonymous):

I am not getting actually sorry... But I am still trying..

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

For A to draw a white ball P(w) = 4/12 For B to draw a white ball P(w0) = 3/11 For C to draw a white ball P(w1) = 2/10 Can you proceed from here?

OpenStudy (anonymous):

after that wat to do..? its just one part..

OpenStudy (anonymous):

add them

OpenStudy (anonymous):

not possible...way off the answer-7/33

OpenStudy (anonymous):

It is not simple as someone can think..

OpenStudy (anonymous):

yup..